### v1

# Topological frequency conversion in rhombohedral multilayer graphene

Abstract
We show that rhombohedral multilayer graphene supports topological frequency conversion, whereby a fraction of electrons transfer energy between two monochromatic light sources at a quantized rate. The pristine nature and gate tunability of these materials, along with a Berry curvature that directly couples to electric fields, make them ideal platforms for the experimental realization of topological frequency conversion. Among the rhombohedral family, we find that Bernal bilayer graphene appears most promising for THz-scale applications due to lower dissipation. We discuss strategies to circumvent cancellations between the two valleys of graphene and to minimize dissipative losses using commensurate frequencies, thus opening a potential pathway for net amplification.
Quantum systems coupled to light exhibit a range of unique emergent phenomena, providing a powerful platform to design and control phases of matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. It was recently appreciated that the paradigm of Floquet engineering can be generalized to multi-chromatic driving, leading to even richer physics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 31]. In this setting, the photon numbers of distinct modes act as synthetic lattice coordinates, opening access to higher-dimensional topological phases that support novel pumping mechanisms. A prime example is topological frequency conversion, whereby energy is transferred between two modes with frequencies $f_{1}$ and $f_{2}$ at a rate given by the power quantum $hf_{1}f_{2}$. In the solid-state realm, Weyl semimetals were recently proposed as a platform to realize this effect, leveraging the sources and sinks of Berry curvature carried by their Weyl points [31].
Here we propose rhombohedral multilayer graphene (RMG) [33, 34] as an alternative platform for realizing topological frequency conversion. Similarly to Weyl semimetals, RMG hosts Weyl nodes, but in a three-dimensional hybrid space formed by the two crystal momenta and the interlayer potential difference. This hybrid space can be dynamically addressed through the in-plane vector potential and out-of-plane electric field induced by the driving modes. Under suitable configurations of incident beams, a fraction of charge carriers in the system converts energy between the modes at the quantized rate $hf_{1}f_{2}$, as illustrated in Fig. 1.
Figure 1: (a) Rhombohedral multilayer graphene can mediate topological frequency conversion between two incident driving modes with frequencies $f_{1}$, $f_{2}$. An interlayer potential $u$ gaps out the low-energy bands of RMG [shown for bilayers in (b) and trilayers in (c)], with a Berry curvature controlled by ${\rm sgn}(u)$. (d) The driving modes couple to electrons via their in-plane vector potential and perpendicular electric field, contained in the 3d vector $\bm{\alpha}$ [Eq. 5]. The Berry curvature ${\bm{\Omega}}$ in $\bm{\alpha}$ space (black arrows) hosts a monopole or Weyl node located at $(-k^{x},-k^{y},0)$ for an electron with wavevector $(k^{x},k^{y})$. Under quasi-periodic driving, the system explores a 2d surface ${\cal S}_{\bm{\alpha}}$, shown for the drive configuration in (a). Frequency conversion power is given by the quantized flux of Berry curvature $\mathbf{\Omega}$ through ${\cal S}_{\bm{\alpha}}$. (e) The intersection between ${\cal S}_{\bm{\alpha}}$ and the $E^{z}=0$ plane defines a momentum area $\cal A_{\bm{\alpha}}$ where electrons contribute to conversion power at a quantized rate $hf_{1}f_{2}L$. Contours are labeled by the ratio of electric fields $E^{z}_{0}/E_{1}$.
The quasi two-dimensional nature of RMG offers key advantages including gate tunability, reduced screening and lack of skin effects. Moreover, the different type of coupling to in-plane and out-of-plane radiation fields leads to a strong sensitivity to the angle and polarization of the incident beams. For THz implementations, Bernal bilayer graphene appears most promising due to its pristine nature and lower dissipation. We finally comment on strategies to circumvent cancellations between the two (time-reversed) valleys of graphene by inducing valley-asymmetric electronic distributions.
Setup. RMG consists of $L$ graphene sheets stacked such that each $A$ sublattice lies above the $B$ sublattice of the layer below—these superimposed sites are coupled with the leading interlayer tunneling term $\gamma_{1}\approx 360$ meV. Near charge neutrality, the wavefunctions of carriers primarily reside in the subspace ($A_{1},B_{L}$) comprising the leftover $A$ sites of the bottom layer and $B$ sites of the top layer. The system can thus be described by an effective two-level Hamiltonian in each of the two valleys [34],
$H_{\xi}(\mathbf{k})=\begin{pmatrix}u/2&-\gamma_{L}\left(\xi k_{x}-ik_{y}\right)^{L}\\ -\gamma_{L}\left(\xi k_{x}+ik_{y}\right)^{L}&-u/2\end{pmatrix},$ (3)
with the wavevector $\mathbf{k}$ measured relative to the hexagonal Brillouin zone (BZ) corner corresponding to valley $\xi=\pm 1$. Here $u$ is an interlayer potential difference and $\gamma_{L}=(\hbar v_{F})^{L}/\gamma_{1}^{L-1}$ where $v_{\rm F}=8.4\times 10^{5}$ m/s denotes the Fermi velocity of graphene. For $u=0$ the low-energy excitations of RMG are chiral fermions with an $L$-th power dispersion [35]. Applying an out-of-plane electric field $E^{z}$ generates a potential $u=d(L-1)eE^{z}$, with $d=0.33$ nm the interlayer distance, opening a gap in the dispersion and controlling the Berry curvature of the low-energy bands (Fig. 1b)—a feature at the heart of our proposal.
Bichromatic driving induces an electric field ${\bf{E}}(t)={\bf{E}}_{1}(t)+{\bf{E}}_{2}(t+\Delta t)$ in the system, where ${\bf{E}}_{j}(t)$ is the monochromatic field from mode $j$ with frequency $f_{j}$, and $\Delta t$ controls the relative phase shift between the modes—relevant when $f_{1}$ and $f_{2}$ are commensurate. Due to the quasi 2D nature of RMG we disregard skin effects and plasmon excitations that become important in 3D implementations such as Weyl semimetals [31]. We further consider THz-scale frequencies, corresponding to wavelengths $\lambda\sim 100\,\rm\mu m$, such that the electric field can be considered uniform in the sample. The in-plane component of the ${\bf{E}}$, ${\bf{E}}^{\parallel}$, couples to electronic crystal momenta via the Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}+e\mathbf{A}^{\parallel}(t)/\hbar$, with ${\bf{A}}^{\parallel}(t)$ the net vector potential induced by ${\bf{E}}^{\parallel}(t)$. In contrast, the out-of-plane component of ${\bf{E}}$, $E^{z}$, couples through the effective dipole moment in the low-energy subspace, Eq. (3). The coupling of radiation thus strongly depends on the incidence angles $\theta_{j}$ and polarization of the drives (Fig. 1a), a unique feature of this setup.
The driven system is conveniently described by the two-level Hamiltonian $H_{\xi}(\mathbf{k},t)={\bf{d}}_{\xi}(\mathbf{k},t)\cdot{\bm{\sigma}}$, with ${\bm{\sigma}}=\left(\sigma^{x},\sigma^{y},\sigma^{z}\right)$ a vector of Pauli matrices acting in the low-energy subspace and
${\bf{d}}_{\xi}(\mathbf{k},t)=\gamma_{L}\left(\text{Re}[\Pi_{\xi}^{L}(\mathbf{k},t)],\text{Im}[\Pi_{\xi}^{L}(\mathbf{k},t)],\alpha^{z}(t)\right).$ (4)
Here $\Pi_{\xi}(\mathbf{k},t)=\xi(k^{x}+\alpha^{x}(t))+i(k^{y}+\alpha^{y}(t))$ and $\bm{\alpha}\equiv\bm{\alpha}_{1}+\bm{\alpha}_{2}$ encode the coupling to electromagnetic fields,
$\bm{\alpha}_{j}(t)=e\left(\frac{A_{j}^{x}(t)}{\hbar},\frac{A_{j}^{y}(t)}{\hbar},\frac{d(L-1)E_{j}^{z}(t)}{2\gamma_{L}}\right).$ (5)
Below, we find it convenient to re-express $\bm{\alpha}_{j}$ (and therefore $\mathbf{d}_{\xi}$) explicitly in terms of the phases $\phi_{j}=\omega_{j}t$, with $\omega_{j}=2\pi f_{j}$ the angular frequency of mode $j$.
Frequency conversion. To see how topological frequency conversion emerges, consider the time-averaged rate of work $P_{j}^{\xi}(\mathbf{k})$ done by mode $j$ on a valence-band electron in valley $\xi$ and wavevector $\mathbf{k}$. Neglecting dissipative effects for now, energy conservation implies $P_{1}^{\xi}(\mathbf{k})=-P_{2}^{\xi}(\mathbf{k})=P_{\rm FC}^{\xi}(\mathbf{k})$, with the frequency conversion power $P_{\rm FC}^{\xi}(\mathbf{k})$ giving the time-averaged energy transfer rate from mode $1$ to mode $2$ mediated by the electron.
When the time dependence of $H_{\xi}(\mathbf{k},t)$ is quasi-adiabatic, $P_{\rm FC}^{\xi}(\mathbf{k})$ is proportional to the number of times the surface spanned by the unit vector $\hat{\mathbf{d}}_{\xi}\equiv{\bf{\mathbf{d}_{\xi}}}/|{\bf{\mathbf{d}_{\xi}}}|$ encloses the origin, $P_{\rm FC}^{\xi}(\mathbf{k})=hf_{1}f_{2}\int_{0}^{2\pi}\frac{d\phi_{1}d\phi_{2}}{4\pi}\hat{{\bf{d}}}_{\xi}\cdot(\partial_{\phi_{1}}\hat{{\bf{d}}}_{\xi}\times\partial_{\phi_{2}}\hat{{\bf{d}}}_{\xi})$ [17, 31, 36]. This result is derived in the Appendix and emerges from the leading-order diabatic correction to the dynamics—along with the fact that, for incommensurate frequencies, the system uniformly explores the two-dimensional parameter space $0\leq\phi_{j}\leq 2\pi$. For commensurate frequencies, this formula is recovered by also averaging $P_{\rm FC}^{\xi}(\mathbf{k})$ over $\Delta t$. Changing variables from $\phi_{j}$ to $\bm{\alpha}$ yields
[…]
In analogy to Weyl semimetals, sources and sinks of Berry curvature in the hybrid three-dimensional space $\bm{\alpha}$ are quantized and confined to topologically-protected gap-closing points, analogous to Weyl nodes  [^1] The term Weyl node normally refers to such nodal points in the parameter space comprising the crystal momenta in 3D reciprocal space. Here we use the term “Weyl node” in a generalized sense to describe gap closings as a function of any three parameters.. Indeed, the RMG Hamiltonian in Eq. (4) hosts a charge-$L$ Weyl node at $\bm{\alpha}_{0}(\mathbf{k})=(-k^{x},-k^{y},0)$,
$\nabla_{\bm{\alpha}}\cdot{\mathbf{\Omega}_{\xi}}(\mathbf{k},\bm{\alpha})=-2\pi\xi L\delta(\bm{\alpha}-\bm{\alpha}_{0}(\mathbf{k})).$ (7)
Electrons therefore become topological frequency converters—mediating an energy transfer at the quantized rate $P_{\rm FC}^{\xi}(\mathbf{k})=-hf_{1}f_{2}\xi L$—if located within the BZ area delineated by the intersection of the surface $\cal S_{\bm{\alpha}}$ and the $\alpha^{z}=0$ plane of the Weyl node, denoted $\cal A_{\bm{\alpha}}$ (see Fig. 1d, e). For electrons outside of $\cal A_{\bm{\alpha}}$, $P_{\rm FC}^{\xi}(\mathbf{k})=0$.
We now consider the dynamics of the many-body state. In the regime of adiabatic driving and slow relaxation relative to the driving frequency, the steady-state population of band $n=0,1$ near valley $\xi$ takes an approximately stationary value, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$  [^2] In a simple relaxation-time approximation, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$ is given by the time-averaged value of $f_{\beta}([\varepsilon_{n}(\mathbf{k},t)-\mu])$ with $f_{\beta}(E)$ the Fermi-Dirac distribution [31]; we expect more genral dissipation mechanisms will lead to qualitatively similar distributions.. The frequency conversion intensity (i.e. the conversion power per unit area) from valley $\xi$ is obtained by weighing the individual contribution $P_{\rm FC}^{\xi}(\mathbf{k})$ of each electron with the corresponding steady-state occupation [31],
${\cal I}_{\rm FC}^{\xi}=hf_{1}f_{2}L\xi\int_{\mathcal{A}_{\alpha}}\frac{{\rm d}^{2}\mathbf{k}}{(2\pi)^{2}}\left(\bar{\rho}_{1}^{\xi}(\mathbf{k})-\bar{\rho}_{0}^{\xi}(\mathbf{k})\right).$ (8)
The boundary of $\mathcal{A}_{\bm{\alpha}}$ contains the wavevectors $\mathbf{k}$ for which the instantaneous gap of $H_{\xi}(\mathbf{k},t)$ closes at some time $t$. States near this boundary thus undergo inter-band Landau-Zener tunneling, which causes significant dissipation through Landau-Zener absorption—whereby electrons pumped to the valence band relax dissipatively, leading to an irreversible loss of the invested energy [31].
To estimate the conversion rate from a single valley at charge neutrality, we set $\bar{\rho}_{0}^{\xi}\approx 1$, $\bar{\rho}_{1}^{\xi}\approx 0$ and approximate $\mathcal{A}_{\bm{\alpha}}$ as a rectangle of sides $2eA_{j}\cos\theta_{j}/\hbar$,
${\cal I}_{\rm FC}^{\xi}\approx{\xi}L\frac{e^{2}E_{1}E_{2}}{2\pi^{3}\hbar}\cos\theta_{1}\cos\theta_{2},$ (9)
[…]
When the chemical potential $\mu_{\xi}$ in a given valley moves away from charge neutrality, we expect the conversion rate $I_{\rm FC}^{\xi}$ to decrease, due to the cancellation between conduction and valence-band carriers (for $\mu_{\xi}>0$) or de-populating the valence band (for $\mu_{\xi}<0$). Because of the opposite contributions from carriers in valleys $\xi=\pm 1$ (a consequence of time-reversal symmetry), a net effect requires a valley imbalance in the number of active frequency converters, either through unequal chemical potentials in each valley, $\mu_{+}\neq\mu_{-}$, or a valley-selective population inversion. We discuss various strategies to induce such imbalances below and in the Appendix.
Numerical simulations. We support our conclusions with numerical simulations of Bernal bilayer graphene subjected to bichromatic radiation, including trigonal warping corrections to the low-energy band structure [39]. Following Ref. [31], we model the system through a master equation
$\partial_{t}\hat{\rho}_{\xi}(\mathbf{k},t)=-\frac{i}{\hbar}[\hat{H}_{\xi}(\mathbf{k},t),\hat{\rho}_{\xi}(\mathbf{k},t)]+{\cal D}_{\xi}(\mathbf{k},t)\circ\hat{\rho}_{\xi}(\mathbf{k},t),$ (10)
where $\hat{H}_{\xi}(\mathbf{k},t)$ and $\hat{\rho}_{\xi}(\mathbf{k},t)$ denote the second-quantized Hamiltonian and density matrix of the system in the Fock space spanned by electronic modes with wavevector $\mathbf{k}$ in valley $\xi$. We model dissipation through ${\cal D}_{\xi}(\mathbf{k},t)=-\frac{1}{\tau}(\hat{\rho}_{\xi}(\mathbf{k},t)-\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t))$, characterizing uniform relaxation on a timescale $\tau$ towards the instantaneous equilibrium state $\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t)=e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}/{\rm Tr}[e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}]$, with $\beta$ the inverse temperature of the bath and $\hat{n}$ the density operator. For simplicity we work at charge neutrality, $\mu_{\xi}=0$; we consider effects of driving in the presence of a Fermi surface (for $\mu_{\xi}\neq 0$) in the Appendix.
Figure 2: Momentum-resolved frequency conversion in Bernal bilayer graphene. We consider two linearly-polarized modes in the configuration of Fig. 1a with $\left(\theta_{j}-\pi/2\right)\approx 0.1^{\circ}$, relaxation time $\tau=500$ ps and charge neutrality, $\mu=0$. Contributions to frequency conversion from valleys $\xi=+1$ (a,d) and $-1$ (b,e) are shown alongside dissipated power (c,f). Frequencies $f_{1}\approx 0.128$ THz and $f_{2}=f_{1}/\sqrt{2}$ [(a)–(c)] or $f_{2}=2f_{1}/3$ [(d)–(f)] illustrate quasi-periodic and periodic driving. The color scales indicate time-averaged conversion power in units of $hf_{1}f_{2}$. The overlaid contour shows the boundary of ${\cal A}_{\alpha}$.
We numerically obtain the steady-state solution to Eq. (10) for a representative grid of $\mathbf{k}$ points and extract the time-averaged power per area transferred into the system from mode $j$ through ${\cal I}_{j}^{\xi}=\int\frac{d^{2}\mathbf{k}}{(2\pi)^{2}}{\cal P}_{j}^{\xi}(\mathbf{k})$, where
${\cal P}_{j}^{\xi}(\mathbf{k})=\lim_{T\rightarrow\infty}\int_{0}^{T}\frac{dt}{T}~{}\partial_{t}\bm{\alpha}_{j}(t)\cdot{\rm Tr}\left[\hat{\rho}_{\xi}(\mathbf{k},t)\nabla_{\bm{\alpha}_{j}}\hat{H}_{\xi}(\mathbf{k},t)\right].$ (11)
This expression can be derived from first principles and, being agnostic to our theoretical analysis, serves as an independent check of its conclusions.
We consider two linearly-polarized modes propagating with incidence angles $\theta_{j}$ (relative to $\hat{\mathbf{z}}$) in the $xz$ and $yz$ planes, as depicted in Fig. 1a. The induced vector potentials are given by $\mathbf{A}_{1}(t)=A_{1}(\cos\theta_{1}\cos[\omega_{1}t],0,\sin\theta_{1}\cos[\omega_{1}t])$ and $\mathbf{A}_{2}(t)=A_{2}(0,-\cos\theta_{2}\cos[\omega_{2}(t{+\Delta t})],\sin\theta_{2}\cos[\omega_{2}(t{+\Delta t})])$, with $\Delta t$ controlling the phase shift between the modes. We consider large angles of incidence $\theta\approx\pi/2$, to compensate for the weaker effective coupling between the radiation fields and $\sigma^{z}$ in Eq. 3. In order to obtain a finite area $\cal A_{\alpha}$ we include a time-independent component of the out-of-plane field, coming from a substrate or a back gate, such that $E^{z}=-\partial_{t}A^{z}+E^{z}_{0}$ (see Fig. 1c). In our simulations, we set $E_{1}=0.1$ V/nm and $E^{z}_{0}\approx 2/3E_{1}$, and tune $E_{2}=rE_{1}$. This electric field scale induces momentum displacements within the region of validity of the low-energy description, $eA_{j}\cos\theta_{j}\lesssim\gamma_{1}/2v_{F}$ [35].
Figure 3: Valley-resolved frequency conversion ${\cal I}_{\rm FC}^{+}$ and dissipation ${\cal I}_{\rm diss}^{+}$ in bilayer graphene as a function of $\Omega=\sqrt{\omega_{1}\omega_{2}}$. The incidence angles $\theta_{j}$ are scaled to keep $\cos\theta_{j}/\omega_{j}$ constant, using Fig. 2 as a reference point. The dashed line $\sim\Omega^{2}$ corresponds to the low-frequency limit of ${\cal I}_{\rm FC}^{+}$, where non-adiabatic corrections can be neglected. We contrast (a) incommensurate and (b) commensurate frequency ratios $r=1/\sqrt{2}$ and $2/3$. For commensurate drives, frequency conversion overpowers dissipative losses in the THz regime.
We first set $\omega_{1}\approx 0.8$ THz and probe different frequency ratios $r=\omega_{2}/\omega_{1}$. Fig. 2 shows the frequency conversion rates ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})=\frac{1}{2}({\cal P}_{\rm 1}^{\pm}(\mathbf{k})-{\cal P}_{2}^{\pm}(\mathbf{k}))$, obtained for $r=1/\sqrt{2}$ (a–b) and $r={2}/{3}$ (d–e), respectively. For the latter case, we compute ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})$ by averaging over $\Delta t$. Consistent with our analysis, electrons located within ${\cal A}_{\bm{\alpha}}$ transfer energy between the modes at the quantized rate $2hf_{1}f_{2}$. Near the boundary, these electrons generate substantial dissipation, ${\cal P}_{\rm diss}(\mathbf{k})=\sum_{\xi}({\cal P}_{\rm 1}^{\xi}(\mathbf{k})+{\cal P}_{2}^{\xi}(\mathbf{k}))$, through the Landau-Zener absorption mechanism [Fig. 2 (c,f)]. Dissipation is strongly reduced for commensurate driving due to the mechanism of Lissajous conversion [31]: Only a few electrons on the boundary of $\mathcal{A}_{\bm{\alpha}}$ are taken to the Weyl node by $e{\bf{A}}(t)/\hbar$ and hence generate Landau-Zener absorption. The anisotropic profile of conversion power along the boundary (panels d-e) arises from the distinct mode polarizations, which leads to a significant $\mathbf{k}$ dependence on the mode most responsible for pumping electrons between bands.
In Fig. 3 we plot the integrated conversion rate from a single valley, ${\cal I}_{\rm FC}^{+}\equiv\frac{1}{2}\left({\cal I}_{1}^{+}-{\cal I}_{2}^{+}\right)$, and the corresponding dissipated power ${\cal I}_{\rm diss}^{+}\equiv({\cal I}_{1}^{+}+{\cal I}_{2}^{+})$ as a function of frequency, with the same values of $r$ as in Fig. 2. To remove geometric factors from the analysis, we scale the incidence angle $\theta_{j}$ such that $\cos\theta_{j}/\omega_{j}$ remains constant, ensuring that the surface $\cal S_{\bm{\alpha}}$ is unchanged (electrons are only moving faster along it). In both cases, ${\cal I}_{\rm FC}^{+}$ initially grows quadratically with frequency, following the scaling of the conversion power quantum $hf_{1}f_{2}L$. This growth is eventually arrested due to the expansion of the boundary region of $\mathcal{A}_{\alpha}$, where electrons exhibit Landau-Zener absorption. This leads to a sub-quadratic scaling and, beyond a critical frequency, a sharp decrease in conversion power. Crucially, for commensurate driving on the THz scale [Fig. 3(b)], ${\cal I}_{\rm FC}^{\xi}$ can vastly outperform dissipative losses with $\tau$ of order $100-1000$ ps. In the Appendix we estimate typical heating rates of order $0.1-1$ K/ps, potentially allowing to perform topological frequency conversion over many THz drive cycles.
Figure 4: Frequency conversion power (a,b) and dissipation (c,d) in valley $\xi=-1$ of rhombohedral trilayer graphene. We consider quasi-periodic (a,c) and periodic (b,d) driving, with parameters identical to Fig. 2. While electrons in the quantized regions contribute $50\%$ more to frequency conversion due to Weyl nodes carrying charge $L=3$, dissipation through Landau-Zener absorption is also enhanced because of the flatter low-energy dispersion.
To demonstrate the extension of topological frequency conversion to other members of the RMG family, we simulate rhombohedral trilayer graphene irradiated by the same bichromatic driving as above (Fig. 4). Our results show a quantized plateau hosting topological frequency conversion at the rate $3hf_{1}f_{2}$ per electron (due to the charge-$3$ Weyl nodes in this system). This gain compared to bilayer graphene is however offset by a larger dissipation rate, due to a flatter low-energy dispersion that enhances the non-adiabatic region—a trend expected to continue with increasing number of layers.
Outlook. We have shown that RMG can mediate energy transfer between electromagnetic modes through the mechanism of topological frequency conversion. The two-dimensional nature of this platform results in an anisotropic coupling to radiation, which in turn leads to a strong dependence of conversion power on the incidence angle and polarization of the drives.
Our results indicate that amplification from a single valley can be achieved with commensurate frequencies in the THz range and relaxation timescales $\tau$ of order $100$ ps. In order to achieve net conversion, however, time-reversal symmetry (which relates the two valleys) must be broken. We speculate that valley-imbalanced populations can be realized either spontaneously through electronic interactions at low temperatures [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], by interfacing with magnetic/spin-orbit coupled materials [52, 49, 53, 54, 55, 56, 57], or through photoexcitations generated by circularly-polarized light [58, 59, 60, 61, 62]. Recent experiments suggest that valley-imbalanced transient states can be long-lived (with valley lifetimes reported in a large range from $1\mu$s to $100$ms in recent bilayer graphene quantum dot experiments with applied magnetic fields [63, 64]). The investigation of such schemes present an interesting opportunity for further studies of frequency conversion phenomena.
Among the RMG family, we find Bernal bilayer graphene to be most promising for achieving THz-scale amplification, due to its steeper low-energy dispersion which minimizes dissipation from Landau-Zener absorption. With low-temperature mean free paths of order $l\sim 10~{}\mu{\rm m}$ [45, 47, 46] and typical Fermi velocities $v\sim 10^{5}$ m/s near charge neutrality, electronic relaxation timescales $\tau\sim l/v\sim 100$ ps appear within experimental reach. Provided that valley-imbalanced populations can be achieved and controlled, our simulations indicate that THz frequency conversion can overpower dissipative loss in this regime and occur before excessive heating sets on.
Finally, topological frequency conversion in RMG requires relatively strong electric fields and precise alignment of the polarization of the incoming radiation. We speculate that both challenges can be addressed by focusing the beams using waveguides and antennas—such structures, along with backgates of suitable geometry providing the required constant electric fields, could become part of a future frequency conversion device.
Acknowledgements.
Acknowledgments. We thank Cyprian Lewandowski and Christopher Yang for insightful discussions and collaborations on related projects. É. L.-H. was supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8682. F.N. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under award DE-SC0019166, the Simons Foundation under award 623768, and the Carlsberg Foundation, grant CF22-0727. G.R. and I.E. are grateful for support from the Simons Foundation and the Institute of Quantum Information and Matter. G.R. is grateful for support from the ARO MURI grant FA9550-22-1-0339. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
Appendix A Topological frequency conversion in slowly-driven two-level systems
In this Appendix we review the physical mechanism of topological frequency conversion, following the approach in Ref. [31], and its application to rhombohedral multilayer graphene. We consider a single valence-band electron at crystal momentum $\mathbf{k}$ and in valley $\xi$, and omit these indices below for notational simplicity.
We start with the general expression for the instantaneous rate of work $P(t)$ done on the system from the driving fields,
$P(t)=\langle\psi(t)|\partial_{t}H(t)|\psi(t)\rangle,$ (12)
where $H(t)=\mathbf{d}(t)\cdot\bm{\sigma}$ is the driven two-level Hamiltonian parametrized by Eq. 4 and $|\psi(t)\rangle$ is the instantaneous state of the electron. As discussed below, $|\psi(t)\rangle$ differs slightly from the instantaneous ground state of $H(t)$ due to diabatic corrections. Expressing the corresponding Bloch vector as ${\bf{v}}(t)=\langle\psi(t)|\bm{\sigma}|\psi(t)\rangle$, we have
$P(t)={\bf{v}}(t)\cdot\partial_{t}{\bf{d}}(t).$ (13)
We now rewrite the time-dependent Hamiltonian Bloch vector ${\bf{d}}(t)$ as a function of the phases of the two modes, $\phi_{1}=\omega_{1}t$ and $\phi_{2}=\omega_{2}[t+\Delta t]$, such that ${\bf{d}}(t)=\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi)$, where $\Delta\phi\equiv\omega_{2}\Delta t$. Then
$\partial_{t}\mathbf{d}(t)=[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi),$ (14)
with the phase difference $\Delta\phi=\omega_{2}\Delta t$. Inserting the above into Eq. (13), we recognize the contribution ${\bf{v}}\cdot\omega_{j}\partial_{\phi_{j}}\mathbf{d}$ as the rate of work done by mode $j$. The time-averaged conversion rate from mode $1$ to mode $2$ is given by $P_{\rm FC}=P_{1}=-P_{2}$ (neglecting dissipation), such that
$P_{\rm FC}=\omega_{1}\lim_{T\to\infty}\int_{0}^{T}\frac{dt}{T}{\bf{v}}(t)\cdot\partial_{\phi_{1}}\mathbf{d}(t).$ (15)
To evaluate the integrand, we note that for quasi-adiabatic driving and slow relaxation, the time-dependent Bloch vector of the system $\mathbf{v}(t)$ can be expressed as
${\bf{v}}=\hat{\mathbf{d}}+\frac{\hbar}{2|\mathbf{d}|}\hat{\mathbf{d}}\times\partial_{t}\hat{\mathbf{d}}+\mathcal{O}\left(\frac{\omega^{2}}{|\mathbf{d}^{2}|}\right),$ (16)
where the second term on the r.h.s. denotes the leading-order diabatic correction and is ultimately responsible for topological frequency conversion. The result Eq. 16 can be derived from a sequence of rotating frame transformations, see e.g. Refs. [65, 36]. We can thus express ${\bf{v}}$ explicitly in terms of the phases $\phi_{1}$ and $\phi_{2}$,
${\bf{v}}=\hat{\mathbf{d}}+\frac{\hbar}{2|\hat{\mathbf{d}}|}\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}}.$ (17)
Plugging this result into Eq. (15) and using $\partial_{\phi_{i}}\mathbf{d}=\hat{\mathbf{d}}\partial_{\phi_{i}}|\mathbf{d}|+|\mathbf{d}|\partial_{\phi_{i}}\hat{\mathbf{d}}$, along with $\hat{\mathbf{d}}\cdot\partial_{\phi_{i}}\hat{\mathbf{d}}=\frac{1}{2}\partial_{\phi_{i}}(\hat{\mathbf{d}}\cdot\hat{\mathbf{d}})=0$, we find
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}=\partial_{\phi_{1}}|\mathbf{d}|+\frac{\hbar}{2}(\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}})\cdot\partial_{\phi_{1}}\hat{\mathbf{d}}$ (18)
where we also used that $(\hat{\mathbf{d}}\times\partial_{i}\hat{\mathbf{d}})\cdot\hat{\mathbf{d}}=0$. Next, we use the identity ${\bf{a}}\cdot({\bf{b}}\times{\bf{c}})={\bf{c}}\cdot({\bf{a}}\times{\bf{b}})$ along with the antisymmetry of the cross product to find
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}=\partial_{\phi_{1}}|\mathbf{d}|{-}\frac{\hbar\omega_{2}}{2}\hat{\mathbf{d}}\cdot(\partial_{{\phi_{1}}}\hat{\mathbf{d}}\times\partial_{{\phi_{2}}}\hat{\mathbf{d}}).$ (19)
We compute $P_{\rm FC}$ by averaging Eq. 19 over time. To this end, we restore explicit time (or phase) dependence and note that ${\bf{v}}(\phi_{1},\phi_{2})\cdot\partial_{\phi_{1}}\mathbf{d}(\phi_{1},\phi_{2})$ is $2\pi$-periodic in its arguments. For incommensurate frequencies, time-averaging is thus equivalent to phase-averaging:
$\lim_{T\to\infty}\int_{0}^{T}\frac{dt}{T}[{\bf{v}}\cdot\partial_{\phi_{1}}\mathbf{d}](\omega_{1}t,\omega_{2}t+\Delta\phi)$
$=\int_{0}^{2\pi}\frac{d^{2}\phi}{4\pi^{2}}[{\bf{v}}\cdot\partial_{\phi_{1}}\mathbf{d}](\phi_{1},\phi_{2}).$ (20)
For commensurate frequencies, $P(\omega_{1}t,\omega_{2}t+\Delta\phi)$ is not identical to phase-averaging becuase the system only explores a one-dimensional, closed contour on the “phase Brillouin zone” $\Phi{\rm BZ}$ spanned by $\phi_{1},\phi_{2}$. However, we recover the incommensurate driving result when also averaging $P_{\rm FC}$ over the phase difference between the two modes $\Delta\phi$. (This procedure was used in our numerical simulations for commensurate driving presented in the main text.)
Finally, using that $\int d^{2}\phi\partial_{\phi_{1}}|\mathbf{d}|=0$ due to periodic boundary conditions, we obtain the final result for the time-averaged (and phase-averaged for commensurate drives) frequency conversion power,
$P_{\rm FC}=-\frac{\hbar\omega_{1}\omega_{2}}{2\pi}\int_{0}^{2\pi}\frac{{\rm d}\phi_{1}\phi_{2}}{4\pi}\hat{\mathbf{d}}\cdot(\partial_{\phi_{1}}\hat{\mathbf{d}}\times\partial_{\phi_{2}}\hat{\mathbf{d}}).$ (21)
Using $\frac{\hbar\omega_{1}\omega_{2}}{2\pi}=hf_{1}f_{2}$, we obtain the result quoted in Eq. 6 of the main text.
Appendix B Scaling of conversion power and dissipation with frequency
Figure 5: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate drives with $r=1/\sqrt{2}$.
Figure 6: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider commensurate drives with $r=2/3$.
In this Appendix we examine how the profile of frequency conversion and dissipation for electrons evolves as a function of frequency. As explained in the main text, we increase the driving frequencies $\omega_{j}$ in tandem with the incidence angles of the modes, $\theta_{j}$, in such a way that $\cos\theta_{j}/\omega_{j}$ remains constant. This ensures that the surface ${\cal S}_{\bm{\alpha}}$ along which the trajectories in $\bm{\alpha}$-space traversed by electrons remain fixed. In our analysis, we keep the ratio of frequencies $r=\omega_{1}/\omega_{2}$ fixed, and vary their geometric mean $\Omega=\sqrt{\omega_{1}\omega_{2}}$.
We contrast incommensurate ($r=1/\sqrt{2}$, shown in Fig. 5) and commensurate ($r=2/3$, shown in Fig. 6) driving. In both cases, the quantized conversion plateau gradually shrinks with $\Omega$, due to the expansion of BZ areas dominated by non-adiabatic effects including, in particular, Landau-Zener absorption. The energy dissipation rate increases roughly linearly with frequency, and therefore becomes progressively less visible when normalized by the quantum of frequency conversion $hf_{1}f_{2}$.
Periodic driving (Fig. 6) possesses two main advantages over quasi-periodic driving (Fig. 5), as discussed in the main text: frequency conversion persists at higher frequencies (panels a–e), and the total dissipated power is smaller (panels f–j). As a result, commensurate driving supports much stronger net conversion at THz-scale frequencies—as demonstrated in Fig. 3 of the main text.
Appendix C Estimate of dissipation rate from nonzero chemical potential
Here we estimate the contribution to dissipation from a nonzero chemical potential $\mu$—that is, from the presence of an equilibrium Fermi surface. We consider the regime where the characteristic relaxation time is much larger than the oscillation period of the driving modes, $\omega\tau\gg 1$ and neglect the effects of the perpendicular displacement field on the band dispersion. We moreover assume $\mu$ to be small, working to leading order in $\mu$. Without loss of generality, we consider the case where $\mu>0$ (negative $\mu$ is related to our results below through particle hole symmetry).
Eq. (10) of the main text implies that the dissipation rate is given by
$\eta_{\rm diss}(t)=\frac{\langle E(t)\rangle-\langle E_{\rm eq}(t)\rangle}{\tau}$ (22)
where
$\langle E(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}(\mathbf{k},t)\hat{H}(\mathbf{k},t)],$ (23)
$\langle E_{\rm eq}(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}_{\rm eq}(\mathbf{k},t)\hat{H}(\mathbf{k},t)].$ (24)
Here $\hat{\cdot}$ indicate operators acting on second-quantized Bloch space.
Since $\hat{H}(\mathbf{k},t)=\hat{H}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$, we have $\hat{\rho}_{\rm eq}(\mathbf{k},t)=\hat{\rho}_{\rm eq}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$. Thus, a shift of integration variable in Eq. (24), $\mathbf{k}\to\mathbf{k}+e{\bf{A}}(t)/\hbar$ shows that $\langle E_{\rm eq}(t)\rangle$ is time-independent and given by
$E_{\rm eq}=\sum_{n}\int\!\!\frac{d^{2}\mathbf{k}^{\prime}}{4\pi^{2}}f_{\rm\beta}(\varepsilon_{n}(\mathbf{k}^{\prime})-\mu)\varepsilon_{n}(\mathbf{k}^{\prime}).$ (25)
where $f_{\beta}(\varepsilon)=1/(1+e^{\beta\varepsilon})$ denotes the Fermi-Dirac distribution and $\varepsilon_{n}(\mathbf{k})$ denote the energies of $H(\mathbf{k},t)$ for $n=0,1$. In the above we used $\mathbf{k}^{\prime}$ for the integration variable, to make the distinction between “stationary” and “comoving” wavevectors clear.
The goal of this Appendix is to compute the contribution to dissipation from intra-band scattering. Inter-band scattering in the form of Landau-Zener absorption is discussed in the main text and arises from nonadiabatic effects. We can hence isolate the dissipation from intra-band scattering by assuming adiabatic dynamics in our treatment [31].
Assuming adiabatic dynamics, and working in the limit $\omega\tau\gg 1$, $\rho(\mathbf{k},t)$ is diagonal in the eigenbasis of $H$, with stationary eigenvalues $p_{n}(\mathbf{k})=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}dtf_{\beta}(\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ [31]. As a result,
$\langle E(t)\rangle=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,p_{n}(\mathbf{k})\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ (26)
The average dissipation from intra-band relaxation is given by
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}(\bar{E}-E_{\rm eq}),$ (27)
with $\bar{E}\equiv\lim_{T\to\infty}\int_{0}^{T}\frac{dt}{T}\langle E(t)\rangle$ denoting the average total energy of all electrons in the system. We compute this quantity by inserting the expression for $p_{n}(\mathbf{k})$ in the definition for $\bar{E}$. Using a change integration variables, $\mathbf{k}\to\mathbf{k}-e{\bf{A}}(t)/\hbar$, we obtain
$\bar{E}=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\bar{p}_{i}(\mathbf{k})\varepsilon_{n}(\mathbf{k})$ (28)
where
$\bar{p}_{n}(\mathbf{k})=\lim_{T_{i}\to\infty}\int_{0}^{T_{1}}\frac{dt_{1}}{T_{1}}\int_{0}^{T_{2}}\frac{dt_{2}}{T_{2}}f_{\beta}\left(\varepsilon_{n}\left(\mathbf{k}+\frac{e}{\hbar}[{\bf{A}}(t_{1})-{\bf{A}}(t_{2})]\right)-\mu\right)$ (29)
Note that $p_{i}(\mathbf{k})$ is obtained from effectively averaging the equilibrium distribution $f_{\beta}(\varepsilon_{i}(\mathbf{k})-\mu)$ over $\mathbf{k}$-points reached by the $\mathbf{k}$-displacements in $\{\frac{e}{\hbar}[{\bf{A}}(t_{1})-{\bf{A}}(t_{2})]|t_{1},t_{2}\in[0,\infty)\}$. Using this result, we obtain
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\varepsilon_{n}(\mathbf{k})[\bar{p}_{n}(\mathbf{k})-f_{\beta}(\varepsilon_{n}-\mu)]$ (30)
To estimate $\bar{\eta}_{\rm mr}$, we approximate $\bar{p}_{0}(\mathbf{k})=1$, and assume the valence band carriers to be uniformly smeared over a disk of radius $\Delta k$ where $\Delta k=3eA_{\parallel}/2\hbar$ with $A_{\parallel}$ denoting the characteristic scale of the in-plane vector potential from the drive (recall that we consider the limit of small $\mu$, where $k_{\rm F}$ in equilibrium is much smaller than $\Delta k$). This results in $\bar{p}_{1}(\mathbf{k})=\frac{k_{\rm F}^{2}}{\Delta k^{2}}\theta(\Delta k-|\mathbf{k}|)$ with $\theta$ denoting the step function. The estimate of $\bar{p}_{1}(\mathbf{k})$ is based on the fact that $\bar{p}(\mathbf{k})$ has maximum at $|\mathbf{k}|=0$, decreases monotonically with $|\mathbf{k}|$, and vanishes for $|\mathbf{k}|\geq 2eA_{\parallel}/\hbar$. We finally assume $f_{\beta}(\varepsilon-\mu)\propto\theta(-(\varepsilon-\mu))$.
Using the above assumptions along with $\varepsilon_{1}(\mathbf{k})={\gamma_{L}}\mathbf{k}^{L}$ and $\varepsilon_{0}(\mathbf{k})=-{\gamma_{L}}\mathbf{k}^{L}$, we obtain
$\bar{\eta}_{\rm mr}\approx\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\frac{\gamma_{L}}{\tau}\mathbf{k}^{L}\left[\theta(\Delta-|\mathbf{k}|)\frac{k_{\rm F}^{2}}{\Delta k^{2}}-\theta(k_{\rm F}-|\mathbf{k}|)\right]$ (31)
Going to polar coordinates, $\mathbf{k}=k(\cos\phi,\sin\phi)$, we obtain
[…]
Evaluating the integral and using $\mu=\gamma_{L}k_{\rm F}^{L}$ we find
$\bar{\eta}_{\rm mr}\approx\frac{\mu k_{\rm F}^{2}}{2\pi\tau(L+2)}[(\Delta k/k_{\rm F})^{L}-1]$ (33)
Using our estimate $\Delta k\approx 3eA_{\parallel}/2\hbar$, we thus obtain
$\bar{\eta}_{\rm mr}\approx\frac{(\mu/\gamma_{L})^{2/L}}{2\pi\tau(L+2)}\left[\gamma_{L}\left(\frac{3eA_{\parallel}}{2\hbar}\right)^{L}-\mu\right].$ (34)
Working to leading order in $\mu$ yields
$\bar{\eta}_{\rm mr}\approx\mu^{2/L}\frac{\gamma_{L}^{2/L-1}}{2\pi\tau(L+2)}\left(\frac{3eA_{\parallel}}{\hbar}\right)^{L}$ (35)
We next use $A_{\parallel}\sim 2A\cos\theta\sim 2E\cos\theta/\omega$ with $A$ and $E$ denoting the characteristic amplitude of the total vector potential and electric field from each drive, respectively, leading to
$\bar{\eta}_{\rm mr}\approx\frac{\mu^{2/L}\gamma_{L}^{1-2/L}}{2\pi\tau(L+2)}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{L}x$ (36)
For $L=2$, this expression is particularly simple:
$\bar{\eta}_{\rm mr}\approx\frac{\mu}{8\pi\tau}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{2}\quad(L=2),$ (37)
We now estimate the dissipation caused by the Fermi surface from this expression, and compare with numerical simulations carried out at non-zero $\mu$ in Figs. 7, 8. Note that Fig. 8 shows the total dissipated power, including both Landau-Zener absorption and momentum relaxation due to the presence of a Fermi surface. We expect that dissipation strength at $\mu=0$ serves as a good proxy for the characteristic rate of Landau-Zener absorption, while the difference in dissipation to the $\mu=0$ value, ${\cal I}_{\rm diss}^{+}(\mu)-{\cal I}_{\rm diss}^{+}(0)$ serves as a reasonable estimate for the momentum-relaxation induced dissipation, $\eta_{\rm mr}$.
[…]
This analysis also predicts that Fermi-surface-induced dissipation stays constant when scaling up the frequencies in conjunction with the incidence angle such that $\cos\theta_{j}/\omega_{j}$ remains fixed, as in Fig. 3 of the main text. We also verify this scaling in Fig. 8. While the $\mu=0$ dissipation increases with frequency due to more prominent Landau-Zener tunneling, the Fermi surface contribution (which turns on with $\mu$) is on the same scale and roughly independent of frequency. Note that we consider incommensurate frequencies; for commensurate frequencies, we expect the dissipation from Landau-Zener absorption to be further suppressed as demonstrated in Fig. 3.
Figure 7: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–d), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels e–h), as a function of the chemical potential $\mu$. Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate frequencies with ratio $r=1/\sqrt{2}$ and parameters identical to those used for Fig. 2a–c in the main text. As the chemical potential is increased, two trends are observed. First, the quantization of the frequency conversion power is gradually lost due to the opposite contribution of the conduction band (panels a–d). Second, the dissipation profile acquires a diffuse contribution from the underlying Fermi surface (panels e–h), which coexists with the sharp “ring of fire” due to Landau-Zener absorption.
Figure 8: Scaling of the valley-resolved frequency conversion intensity ${\cal I}^{+}_{\rm FC}$ (left) and the corresponding dissipated power ${\cal I}^{+}_{\rm diss}$ (right), as a function of the chemical potential $\mu$ and different frequencies $\Omega=\sqrt{\omega_{1}\omega_{2}}$ as in Figs. 2 and 7. The data for the lowest $\Omega$ are obtained by integrating the power distributions shown in Fig. 7 (a-d) and (e-h), respectively.
Appendix D Estimate of the heating rate from driving
[…]